# Data-efficient surrogate modeling of spectral functions using Gaussian processes: An application to the $t^{\prime}-t′′t^{\prime\prime}$-$J$ model

### 30 Mar 2026 v2

@@ Abstract
Spectral functions encode key many-body information but are costly to compute with high fidelity. Machine-learning surrogates have emerged as a powerful alternative, yet many approaches require large training datasets. We develop a data-efficient surrogate for spectral functions using the $t^{\prime}-t′′t^{\prime\prime}$-$J$ model, which describes the motion of a hole in a quantum antiferromagnet. Using $\sim$ 105 self-consistent Born approximation-based spectra from Lee, Carbone and Yin (Phys. Rev. B 107, 205132 (2023)), we train a deep-kernel Gaussian process surrogate model with sparse variational inference (DKL-SVGP) using only 10% of the available training spectra. We benchmark against feed-forward neural networks (FFNN) trained on the same reduced subset and on the full dataset. The proposed DKL-SVGP model consistently outperforms the reduced-data FFNN and, despite using only 10% of the training spectra, achieves spectrum-wise errors within the same order-of-magnitude as the full-data FFNN baseline. Worst-tail diagnostics show improved fidelity on difficult spectra, while peak-level analysis indicates that DKL-SVGP recovers dominant peak heights with comparable accuracy and improves peak-location agreement under a matched-peak evaluation that mitigates rare peak-swapping cases. Overall, these results highlight GP-based surrogates as a competitive and data-efficient approach for spectral-function prediction in scarce-data regimes.
@@ 1 Introduction
Spectral functions sit at the heart of the modern condensed matter physics. They provide valuable information on one- and many-body excitation spectra, such as quasiparticle dispersions, lifetimes, and emergent low energy excitations. In addition, they are indispensable for understanding and interpreting transport and thermodynamic measurements and are the direct target of spectroscopic probes such as ARPES, x-ray, and neutron scattering Damascelli et al. (2003); Boschini et al. (2024); Ament et al. (2011). At the same time, obtaining high-fidelity spectral functions often remains computationally demanding.
@@
This difficulty spans across a wide range of settings: impurity solvers, such as numerical renormalization group (NRG) and quantum Monte Carlo (QMC) methods, as well as diagrammatic and exact diagonalization techniques can be accurate but become costly when scanning multi-parameter Hamiltonians or when repeated evaluations are needed for inference and parameter exploration. This computational bottleneck motivates machine-learning surrogate models  that learn either the forward map (Hamiltonian parameters $\rightarrow$ spectral function $A(\omega)$) or the inverse map ($A(\omega)$ $\rightarrow$ parameters), with the goal of accelerating parameter exploration and enabling data-driven inference Arsenault et al. (2014); Sturm et al. (2021); Lee et al. (2023). In practice, the learned map is often highly structured: a small number of Hamiltonian parameters can induce nontrivial changes in peak locations, widths, and spectral weight redistribution across the full energy window. This makes spectral surrogates an appealing testbed for data-efficient regression methods that can exploit smoothness while maintaining good accuracy with limited training data and can generalize across the underlying parameter space.
@@
A number of recent works have demonstrated that modern machine learning (ML) can accurately interpolate spectral data when sufficiently large training sets are available. The pioneering work at this front was done by Arsenault et al. in 2014 Arsenault et al. (2014) where they applied various machine learned representations to calculate spectral function for the Anderson Impurity model (AIM), a paradigmatic model of quantum impurity. Building on this, Sturm et al. Sturm et al. (2021) in 2021 built a high-fidelity large NRG database ($\sim$5×105 spectra) and compared deep neural network (DNN)-based model to kernel ridge regression approach, with DNN generally superior at scale. Beyond forward prediction, Miles et al. in 2021 used variational autoencoders to learn interpretable latent variables from single-impurity Anderson model (SIAM) spectra Miles et al. (2021). More recently, Lee, Carbone, and Yin (LCY) (referred to as LCY hereafter) in 2023 generated $\sim 1.3\times 10^{5}$ self-consistent Born approximation (SCBA) spectra for the $t^{\prime}t′′t^{\prime\prime}$$J$ model describing the motion of a single hole in a quantum antiferromagnet Lee et al. (2023). They showed that simple feed-forward neural networks can accurately learn both the forward map and inverse parameter inference, yielding substantial speedups relative to direct SCBA evaluation Lee et al. (2023). Related progress has also appeared in ARPES-oriented ML workflows and in data-driven treatments of spectral reconstruction tasks (e.g., analytic continuation, denoising etc), where training data can be limited and robustness is essential Ekahana et al. (2023); Na et al. (2025); Bian et al. (2024); Kim et al. (2021).
@@
At the same time, two practical issues repeatedly arise when moving from proof-of-principle demonstrations toward workflows for scientific discovery. First, many spectral surrogates still require large training sets, whereas in many applications high-quality spectra are expensive to simulate and only a modest number of computer model-based solver calls are feasible under tight compute budgets. Second, point predictions can be difficult to use in downstream tasks such as active learning, experimental design, or inverse problems, where it is helpful to know when the surrogate is operating within its training regime versus extrapolating Donatelli et al. (2025); related surrogate and inference workflows that incorporate reliability and error-awareness are now routine in other computational science settings, including probabilistic projections Jantre et al. (2024) and computer model system-identification Jantre et al. (2025). In spectroscopy ML, lightweight reliability indicators can be useful, especially when out-of-distribution inputs and systematic errors are possible, and they can help contextualize model predictions Ghose et al. (2023); Carbone et al. (2024).
@@
These considerations naturally motivate Gaussian-process (GP) surrogates for spectral-function modeling in low-dimensional Hamiltonian parameterizations. GP regression provides a flexible nonparametric surrogate that encodes smoothness through its kernel and is well known to be effective in small-data regimes Rasmussen and Williams (2005). From a physics perspective, GP priors can be viewed as enforcing controlled smoothness assumptions on the parameter-to-spectrum map, while allowing the data to determine the relevant correlation lengthscales. Moreover, GP models connect naturally to sequential design and active-learning workflows Santner et al. (2003); Cohn et al. (1994), and GP-based calibration frameworks provide a principled way to relate simulator outputs to experimental data in the presence of model discrepancy Kennedy and O’Hagan (2001). The main historical obstacle has been scalability: exact GP regression scales cubically with the number of training data samples. However, sparse variational GP methods using inducing variables and stochastic optimization make GP regression practical at the $10^{4}$$10^{5}$ scale Titsias (2009); Hensman et al. (2013). Deep kernel learning (DKL) combines neural feature extraction with GP-based predictions by learning a data-adapted kernel, and stochastic variational formulations enable DKL to scale to large data while retaining GP approaches’ nonparametric modeling advantages Wilson et al. (2016).
@@
In this work, we develop a data-efficient surrogate for the LCY’s SCBA database of the $t^{\prime}t′′t^{\prime\prime}$$J$ model. We focus on the forward task of predicting the density of states $A(\omega)$ from the three-dimensional Hamiltonian parameter vector $(t^{\prime},t^{\prime\prime},J)$ (with $t$ setting the overall energy scale as in LCY Lee et al. (2023)). Using only a uniformly random 10% subset of the available $\sim 10^{5}$ spectra, we train a deep-kernel stochastic variational Gaussian process (DKL-SVGP) surrogate model and benchmark against feed-forward neural networks trained on the same reduced budget as well as on the full training dataset. We find that our GP surrogate consistently outperforms the FFNN trained on the same reduced dataset and, despite using only 10% of the training spectra, attains spectrum-wise errors of the same order-of-magnitude as the full-data FFNN baseline, while maintaining comparable peak-height accuracy. Peak locations are typically accurate, but we observe rare large deviations motivating an explicit peak-matching diagnostic protocol that substantially reduces these outliers and improves peak-location fidelity. Overall, these results show that our GP-based surrogate offers a compelling accuracy-data tradeoff for spectral-function modeling in scarce-data regimes and provides a natural foundation for active learning and inverse modeling workflows.
@@ 2.1 Overview of the Physical System
We study the spectral function of the $t^{\prime}-t′′t^{\prime\prime}$-$J$ model Yin et al. (1998); Manousakis (2007b), which is an improved description of the tJt-J model, originally proposed to describe the motion of a single hole in a quantum antiferromagnet Marsiglio et al. (1991); Martinez and Horsch (1991); Liu and Manousakis (1992); Yin and Gong (1997); Manousakis (2007a). For completeness, we write the Hamiltonian in Eq. 1 and briefly summarize how the density of states (DOS) data used for training are generated.
@@
$A(\omega)$ is computed with the non-crossing self-consistent Born approximation (SCBA) Schmitt-Rink et al. (1988); Liu and Manousakis (1991), which iteratively evaluates the hole Green’s function dressed by magnon excitations until convergence. In practice, SCBA requires dense sampling in both the hole momentum $\mathbf{k}$ and magnon momentum $\mathbf{q}$; the dataset we use corresponds to converged calculations on the grids reported in LCY Lee et al. (2023).
@@ 2.2 Machine Learning Models
where $T$ represents the evaluation data. We tune hyperparameters on the validation set using grid search over the network architecture (netnet), batch size (bs), learning rate (lr), and learning-rate scheduler (lrs). For the reduced training set (size $\sim 10^{4}$), the best-performing FFNN uses net=(3, 32, 64, 128, 256, 301)net=(3,\,32,\,64,\,128,\,256,\,301), $\mathrm{bs}=128$, $\mathrm{lr}=5\times 10^{-3}$, and $\mathrm{lrs}=0.5$. For reference and to compare against a high-capacity baseline trained on the full dataset, we also consider the FFNN architecture used in LCY Lee et al. (2023), net=(3, 170, 340, 510, 680, 850, 1020, 301)net=(3,\,170,\,340,\,510,\,680,\,850,\,1020,\,301) with $\mathrm{bs}=1024$, $\mathrm{lr}=10^{-3}$, and $\mathrm{lrs}=0.5$.
@@ A GP places a prior directly on the unknown function,
A key practical issue is scalability: exact GP training scales as $\mathcal{O}(N^{3})$ in the number of training points $N$, and becomes prohibitive as the number of training pairs $N_{\text{pairs}}=N_{\text{spec}}\times P$ grows. We therefore use a sparse variational formulation with inducing variables, which makes GP learning practical at our dataset size Titsias (2009); Hensman et al. (2013). To improve expressivity beyond a stationary kernel on raw inputs, we adopt deep kernel learning (DKL): a lightweight neural network feature map $\phi_{\theta}(\cdot)$ is learned jointly with the GP so that similarity is measured in a task-adapted representation rather than directly in $\mathbf{x},\omega]$ space Wilson et al. (2016). Concretely, we use an end-to-end DKL-SVGP surrogate model with an implicit kernel of the form $k(\mathbf{s},\mathbf{s}^{\prime})=k_{\mathrm{base}}(\phi_{\theta}(\mathbf{s}),\phi_{\theta}(\mathbf{s}^{\prime}))$, trained with minibatch stochastic optimization. Figure 1 summarizes the model architecture and data flow.
@@ 3 Results and Discussion
Table 1: Per-spectrum error (row-RMSE) across data splits. Values are mean and median (in parentheses) over spectra in each data split and are shown as $10^{-3}\!\times\!\text{row-RMSE}$ (lower is better). “FFNN (full)” is the full training data baseline from LCY; “FFNN Subset” and “DKL-SVGP” are trained on the same uniformly randomly selected $10\%$ subset of full training data from LCY Lee et al. (2023).
@@
First, under the same scarce-data training budget (a uniformly random $10\%$ subset of the LCY training set), the DKL-SVGP surrogate is substantially more accurate than the corresponding reduced-data FFNN baseline. On the test split, DKL-SVGP reduces the mean row-RMSE from $7.831\times 10^{-3}$ (FFNN Subset) to $4.544\times 10^{-3}$, i.e., a $\approx 42\%$ reduction, while the median row-RMSE drops from $6.119\times 10^{-3}$ to $3.307\times 10^{-3}$ (a $\approx 46\%$ reduction). A similar improvement is observed in the normalized metric: the mean test NRMSE decreases from $59.449\times 10^{-3}$ to $33.685\times 10^{-3}$ (a $\approx 43\%$ reduction) and the median test NRMSE reduces from $47.755\times 10^{-3}$ to $25.571\times 10^{-3}$ (a $\approx 46\%$ reduction). This indicates that the gain is not driven solely by a few high-amplitude spectra but persists after per-spectrum normalization. This behavior is aligned with the qualitative intuition emphasized in LCY’s forward-problem analysis: interpolation quality is most meaningfully judged by whether the surrogate preserves spectral morphology–peak locations, widths, and heights–even on the hardest examples in the test set. In that spirit, our worst-tail diagnostic (Fig. 2) mirrors the LCY presentation of “worst percentiles” and provides a stringent stress test focused on difficult spectra rather than average-case performance Lee et al. (2023). In these worst-tail panels, the DKL-SVGP predictions generally track the dominant peak structure and capture secondary features more reliably than the reduced-data FFNN, consistent with the quantitative advantage in Tables 12.
@@
Finally, many downstream uses of spectral surrogates depend on peak-level physics; for example, peak positions encode information about excitation energies and band dispersion, whereas peak heights are related to the quasiparticle weights. Therefore, we complement spectrum-wise errors with peak diagnostics (Fig. 3). Scatter plots of the dominant peak height $A_{\max}$ and corresponding location $\omega_{\max}$ assess how well each surrogate preserves salient low-energy structure, in the same spirit as LCY’s emphasis on reproducing peak positions and heights when discussing challenging spectra Lee et al. (2023). To this end, we compare predicted versus ground-truth dominant peak height $A_{\max}$ (top row) and peak location $\omega_{\max}$ (bottom row), using the matched-peak definition in Section 2.3 to reduce sensitivity to rare cases of dominant/secondary (top-$K$ matching with $K=5$). Under this diagnostic, peak-height accuracy is broadly comparable between the two reduced-data models: the FFNN trained on a 10% subset attains RMSE $\approx 0.053$ and MAE $\approx 0.026$, while DKL-SVGP yields RMSE $\approx 0.056$ and MAE $\approx 0.026$ (both with strong correlations $\gtrsim 0.98$). Peak-location accuracy is better for DKL-SVGP: it attains RMSE $\approx 0.021$ compared to $\approx 0.028$ for the reduced-data FFNN, MAE $\approx 0.010$ versus $\approx 0.016$, with correlations $\gtrsim 0.99$ for both. For completeness, we report diagnostic results under strict argmax definition and related outlier diagnostics in Appendix C, where the largest peak-location deviations are shown to arise primarily from occasional peak swapping rather than a complete loss of spectral fidelity.

### 07 May 2026 v3

@@ Abstract
Spectral functions encode key many-body information but are costly to compute with high fidelity. Machine-learning surrogates have emerged as a powerful alternative, yet many approaches require large training datasets. We develop a data-efficient surrogate for spectral functions using the $t^{\prime}-t′′t^{\prime\prime}$-$J$ model, which describes the motion of a hole in a quantum antiferromagnet. Using $\sim$ 105 self-consistent Born approximation-based spectra from Lee, Carbone and Yin (Phys. Rev. B 107, 205132 (2023)), we train a deep-kernel Gaussian process surrogate model with sparse variational inference (DKL-SVGP) using only 10% of the available training spectra. We benchmark against feed-forward neural networks (FFNN) trained on the same reduced subset and on the full dataset. The proposed DKL-SVGP model consistently outperforms the reduced-data FFNN and, despite using only 10% of the training spectra, achieves spectrum-wise errors within the same order-of-magnitude as the full-data FFNN baseline. Worst-tail diagnostics show improved fidelity on difficult spectra, while peak-level analysis indicates that DKL-SVGP recovers dominant peak heights with comparable accuracy and improves peak-location agreement under a matched-peak evaluation that mitigates rare peak-swapping cases. Overall, these results highlight GP-based surrogates as a competitive and data-efficient approach for spectral-function prediction in scarce-data regimes. The DKL-SVGP model is implemented in the open source software Python and is available at: https://github.com/jsanket12/Spectral_fn/.
@@ 1 Introduction
Spectral functions sit at the heart of the modern condensed matter physics. They provide valuable information on one- and many-body excitation spectra, such as quasiparticle dispersions, lifetimes, and emergent low energy excitations. In addition, they are indispensable for understanding and interpreting transport and thermodynamic measurements and are the direct target of spectroscopic probes such as ARPES, x-ray, and neutron scattering (Damascelli et al., 2003; Boschini et al., 2024; Ament et al., 2011). At the same time, obtaining high-fidelity spectral functions often remains computationally demanding.
@@
This difficulty spans across a wide range of settings: impurity solvers, such as numerical renormalization group (NRG) and quantum Monte Carlo (QMC) methods, as well as diagrammatic and exact diagonalization techniques can be accurate but become costly when scanning multi-parameter Hamiltonians or when repeated evaluations are needed for inference and parameter exploration. This computational bottleneck motivates machine-learning surrogate models  that learn either the forward map (Hamiltonian parameters $\rightarrow$ spectral function $A(\omega)$) or the inverse map ($A(\omega)$ $\rightarrow$ parameters), with the goal of accelerating parameter exploration and enabling data-driven inference (Arsenault et al., 2014; Sturm et al., 2021; Lee et al., 2023). In practice, the learned map is often highly structured: a small number of Hamiltonian parameters can induce nontrivial changes in peak locations, widths, and spectral weight redistribution across the full energy window. This makes spectral surrogates an appealing testbed for data-efficient regression methods that can exploit smoothness while maintaining good accuracy with limited training data and can generalize across the underlying parameter space.
@@
A number of recent works have demonstrated that modern machine learning (ML) can accurately interpolate spectral data when sufficiently large training sets are available. The pioneering work at this front was done by Arsenault et al. (2014) where they applied various machine learned representations to calculate spectral function for the Anderson Impurity model (AIM), a paradigmatic model of quantum impurity. Building on this, Sturm et al. (2021) built a high-fidelity large NRG database ($\sim$5×105 spectra) and compared deep neural network (DNN)-based model to kernel ridge regression approach, with DNN generally superior at scale. Beyond forward prediction, Miles et al. (2021) used variational autoencoders to learn interpretable latent variables from single-impurity Anderson model (SIAM) spectra. More recently, Lee, Carbone, and Yin (LCY) (referred to as LCY hereafter) in 2023 generated $\sim 1.3\times 10^{5}$ self-consistent Born approximation (SCBA) spectra for the $t^{\prime}t′′t^{\prime\prime}$$J$ model describing the motion of a single hole in a quantum antiferromagnet (Lee et al., 2023). They showed that simple feed-forward neural networks can accurately learn both the forward map and inverse parameter inference, yielding substantial speedups relative to direct SCBA evaluation (Lee et al., 2023). Related progress has also appeared in ARPES-oriented ML workflows and in data-driven treatments of spectral reconstruction tasks (e.g., analytic continuation, denoising etc), where training data can be limited and robustness is essential (Ekahana et al., 2023; Na et al., 2025; Bian et al., 2024; Kim et al., 2021).
@@
At the same time, two practical issues repeatedly arise when moving from proof-of-principle demonstrations toward workflows for scientific discovery. First, many spectral surrogates still require large training sets, whereas in many applications high-quality spectra are expensive to simulate and only a modest number of computer model-based solver calls are feasible under tight compute budgets. Second, point predictions can be difficult to use in downstream tasks such as active learning, experimental design, or inverse problems, where it is helpful to know when the surrogate is operating within its training regime versus extrapolating (Donatelli et al., 2025); related surrogate and inference workflows that incorporate reliability and error-awareness are now routine in other computational science settings, including probabilistic projections (Jantre et al., 2024) and computer model system-identification (Jantre et al., 2025). In spectroscopy ML, lightweight reliability indicators can be useful, especially when out-of-distribution inputs and systematic errors are possible, and they can help contextualize model predictions (Ghose et al., 2023; Carbone et al., 2024).
@@
These considerations naturally motivate Gaussian-process (GP) surrogates for spectral-function modeling in low-dimensional Hamiltonian parameterizations. GP regression provides a flexible nonparametric surrogate that encodes smoothness through its kernel and is well known to be effective in small-data regimes (Rasmussen and Williams, 2005). From a physics perspective, GP priors can be viewed as enforcing controlled smoothness assumptions on the parameter-to-spectrum map, while allowing the data to determine the relevant correlation lengthscales. Moreover, GP models connect naturally to sequential design and active-learning workflows (Santner et al., 2003; Cohn et al., 1994), and GP-based calibration frameworks provide a principled way to relate simulator outputs to experimental data in the presence of model discrepancy (Kennedy and O’Hagan, 2001). The main historical obstacle has been scalability: exact GP regression scales cubically with the number of training data samples. However, sparse variational GP methods using inducing variables and stochastic optimization make GP regression practical at the $10^{4}$$10^{5}$ scale (Titsias, 2009; Hensman et al., 2013). Deep kernel learning (DKL) combines neural feature extraction with GP-based predictions by learning a data-adapted kernel, and stochastic variational formulations enable DKL to scale to large data while retaining GP approaches’ nonparametric modeling advantages (Wilson et al., 2016).
@@
In this work, we develop a data-efficient surrogate for the LCY’s SCBA database of the $t^{\prime}t′′t^{\prime\prime}$$J$ model. We focus on the forward task of predicting the density of states $A(\omega)$ from the three-dimensional Hamiltonian parameter vector $(t^{\prime},t^{\prime\prime},J)$ (with $t$ setting the overall energy scale as in LCY). Using only a uniformly random 10% subset of the available $\sim 10^{5}$ spectra, we train a deep-kernel stochastic variational Gaussian process (DKL-SVGP) surrogate model and benchmark against feed-forward neural networks trained on the same reduced budget as well as on the full training dataset. We find that our GP surrogate consistently outperforms the FFNN trained on the same reduced dataset and, despite using only 10% of the training spectra, attains spectrum-wise errors of the same order-of-magnitude as the full-data FFNN baseline, while maintaining comparable peak-height accuracy. Peak locations are typically accurate, but we observe rare large deviations motivating an explicit peak-matching diagnostic protocol that substantially reduces these outliers and improves peak-location fidelity. Overall, these results show that our GP-based surrogate offers a compelling accuracy-data tradeoff for spectral-function modeling in scarce-data regimes and provides a natural foundation for active learning and inverse modeling workflows.
@@ 2.1 Overview of the Physical System
We study the spectral function of the $t^{\prime}-t′′t^{\prime\prime}$-$J$ model (Yin et al., 1998; Manousakis, 2007b), which is an improved description of the tJt-J model, originally proposed to describe the motion of a single hole in a quantum antiferromagnet (Marsiglio et al., 1991; Martinez and Horsch, 1991; Liu and Manousakis, 1992; Yin and Gong, 1997; Manousakis, 2007a). For completeness, we write the Hamiltonian in Eq. 1 and briefly summarize how the density of states (DOS) data used for training are generated.
@@
$A(\omega)$ is computed with the non-crossing self-consistent Born approximation (SCBA) (Schmitt-Rink et al., 1988; Liu and Manousakis, 1991), which iteratively evaluates the hole Green’s function dressed by magnon excitations until convergence. In practice, SCBA requires dense sampling in both the hole momentum $\mathbf{k}$ and magnon momentum $\mathbf{q}$; the dataset we use corresponds to converged calculations on the grids reported in LCY.
@@ 2.2 Machine Learning Models
where $T$ represents the evaluation data. We tune hyperparameters on the validation set using grid search over the network architecture (netnet), batch size (bs), learning rate (lr), and learning-rate scheduler (lrs). For the reduced training set (size $\sim 10^{4}$), the best-performing FFNN uses net=(3, 32, 64, 128, 256, 301)net=(3,\,32,\,64,\,128,\,256,\,301), $\mathrm{bs}=128$, $\mathrm{lr}=5\times 10^{-3}$, and $\mathrm{lrs}=0.5$. For reference and to compare against a high-capacity baseline trained on the full dataset, we also consider the FFNN architecture used in LCY, net=(3, 170, 340, 510, 680, 850, 1020, 301)net=(3,\,170,\,340,\,510,\,680,\,850,\,1020,\,301) with $\mathrm{bs}=1024$, $\mathrm{lr}=10^{-3}$, and $\mathrm{lrs}=0.5$.
@@ A GP places a prior directly on the unknown function,
A key practical issue is scalability: exact GP training scales as $\mathcal{O}(N^{3})$ in the number of training points $N$, and becomes prohibitive as the number of training pairs $N_{\text{pairs}}=N_{\text{spec}}\times P$ grows. We therefore use a sparse variational formulation with inducing variables, which makes GP learning practical at our dataset size (Titsias, 2009; Hensman et al., 2013). To improve expressivity beyond a stationary kernel on raw inputs, we adopt deep kernel learning (DKL): a lightweight neural network feature map $\phi_{\theta}(\cdot)$ is learned jointly with the GP so that similarity is measured in a task-adapted representation rather than directly in $\mathbf{x},\omega]$ space (Wilson et al., 2016). Concretely, we use an end-to-end DKL-SVGP surrogate model with an implicit kernel of the form $k(\mathbf{s},\mathbf{s}^{\prime})=k_{\mathrm{base}}(\phi_{\theta}(\mathbf{s}),\phi_{\theta}(\mathbf{s}^{\prime}))$, trained with minibatch stochastic optimization. Figure 1 summarizes the model architecture and data flow.
@@ 3 Results and Discussion
Table 1: Per-spectrum error (row-RMSE) across data splits. Values are mean and median (in parentheses) over spectra in each data split and are shown as $10^{-3}\!\times\!\text{row-RMSE}$ (lower is better). “FFNN (full)” is the full training data baseline from LCY; “FFNN Subset” and “DKL-SVGP” are trained on the same uniformly randomly selected $10\%$ subset of full training data from LCY.
@@
First, under the same scarce-data training budget (a uniformly random $10\%$ subset of the LCY training set), the DKL-SVGP surrogate is substantially more accurate than the corresponding reduced-data FFNN baseline. On the test split, DKL-SVGP reduces the mean row-RMSE from $7.831\times 10^{-3}$ (FFNN Subset) to $4.544\times 10^{-3}$, i.e., a $\approx 42\%$ reduction, while the median row-RMSE drops from $6.119\times 10^{-3}$ to $3.307\times 10^{-3}$ (a $\approx 46\%$ reduction). A similar improvement is observed in the normalized metric: the mean test NRMSE decreases from $59.449\times 10^{-3}$ to $33.685\times 10^{-3}$ (a $\approx 43\%$ reduction) and the median test NRMSE reduces from $47.755\times 10^{-3}$ to $25.571\times 10^{-3}$ (a $\approx 46\%$ reduction). This indicates that the gain is not driven solely by a few high-amplitude spectra but persists after per-spectrum normalization. This behavior is aligned with the qualitative intuition emphasized in LCY’s forward-problem analysis: interpolation quality is most meaningfully judged by whether the surrogate preserves spectral morphology–peak locations, widths, and heights–even on the hardest examples in the test set. In that spirit, our worst-tail diagnostic (Fig. 2) mirrors the LCY presentation of “worst percentiles” and provides a stringent stress test focused on difficult spectra rather than average-case performance (Lee et al., 2023). In these worst-tail panels, the DKL-SVGP predictions generally track the dominant peak structure and capture secondary features more reliably than the reduced-data FFNN, consistent with the quantitative advantage in Tables 12.
@@
Finally, many downstream uses of spectral surrogates depend on peak-level physics; for example, peak positions encode information about excitation energies and band dispersion, whereas peak heights are related to the quasiparticle weights. Therefore, we complement spectrum-wise errors with peak diagnostics (Fig. 3). Scatter plots of the dominant peak height $A_{\max}$ and corresponding location $\omega_{\max}$ assess how well each surrogate preserves salient low-energy structure, in the same spirit as LCY’s emphasis on reproducing peak positions and heights when discussing challenging spectra (Lee et al., 2023). To this end, we compare predicted versus ground-truth dominant peak height $A_{\max}$ (top row) and peak location $\omega_{\max}$ (bottom row), using the matched-peak definition in Section 2.3 to reduce sensitivity to rare cases of dominant/secondary (top-$K$ matching with $K=5$). Under this diagnostic, peak-height accuracy is broadly comparable between the two reduced-data models: the FFNN trained on a 10% subset attains RMSE $\approx 0.053$ and MAE $\approx 0.026$, while DKL-SVGP yields RMSE $\approx 0.056$ and MAE $\approx 0.026$ (both with strong correlations $\gtrsim 0.98$). Peak-location accuracy is better for DKL-SVGP: it attains RMSE $\approx 0.021$ compared to $\approx 0.028$ for the reduced-data FFNN, MAE $\approx 0.010$ versus $\approx 0.016$, with correlations $\gtrsim 0.99$ for both. For completeness, we report diagnostic results under strict argmax definition and related outlier diagnostics in Appendix C, where the largest peak-location deviations are shown to arise primarily from occasional peak swapping rather than a complete loss of spectral fidelity.